# Zero-Shot Image Denoising for High-Resolution Electron Microscopy

### v1

@@ I Introduction
High-resolution electron microscopy (HREM) imaging [1, 2, 3] is an indispensable tool in the fields of materials science and nanotechnology. HREM enables direct visualization of structures at the atomic level through interactions between the sample and high-energy electrons.
@@
However, HREM is inevitably susceptible to noise due to inherent properties of electron beams and detection process etc. For example, the low-dose conditions are often applied to imaging electron-beam sensitive materials to minimize the damage from the electrons. [4, 5]. Capuring the dynamic events at kilohertz frame rates with direct electronic detection systems [6, 7], the image is affected by severe shot noise due to the shortened exposure time. Improving signal-to-noise (SNR) is critical for HREM to enhance image quality and facilitate accurate information extraction.
@@
Recently, data-driven methods based on deep learning (DL) [8, 9, 10] have obtained favorable performance compared to conventional methods in image denoising. However, applying supervised DL denoising methods to electron microscopy (EM) images is challenging due to the lack of paired noisy-clean image datasets. In the absence of ground-truth images, several self-supervised image denoising methods[11, 12, 13, 14, 15, 16, 17, 18, 19] have been proposed. Some works [12, 11, 13, 17] utilize blind-spot networks (BSNs) to prevent identical mappings in self-supervised learning. BSN aims to eliminate the influence of each pixel on the corresponding output pixel to satisfy the $\mathcal{J}$-invariance [11] theory. Noise2Void [12] and Noise2Self [11] employ a masked-based strategy while subsequent works [14, 13, 20, 17] design tailored networks to build BSN. The scarcity of HREM data poses challenges for BSN-based methods, and their performance tends to degrade when trained with limited data [15]. Huang et al. proposes Neighbor2Neighbor (NB2NB) [19] to generate paired noisy images from subsampling paired neighbor pixels for self-supervised training. While NB2NB and its variants [21, 22] fail to address the intensity gap issue between neighbor pixels, resulting in relatively poor performance in low-SNR scenarios. Some works [18, 23] introduce explicit noise modeling to generate training paired noisy images by adding synthetic noise. However, the noise distribution of real HREM is unknown and challenging to estimate due to the extremely low signal-to-noise ratio in HREM images.
@@
In this paper, we propose an efficient zero-shot self-supervised denoising framework for HREM images named Noise2SR. We propose to train a denoising network with paired noisy images with different resolutions, which is conducted via super-resolution (SR) strategy. Inspired by NB2NB [19], we introduce the Random Sub-sampler module to generate sub-sampled noisy images that form a noisy pair with the original noisy image. Unlike NB2NB, we utilize paired noisy images with different resolutions for training. We indicated that the sub-sampled noisy image and the original noisy image have a consistent underlying clean image. Meanwhile, we provide theoretical proof that the Noise2SR training scheme is statically equivalent to using a clean image for supervision. Combined with the SR-based training strategy, we address the coordinate mismatch and intensity gap issues present in paired neighbor pixels in NB2NB. The proposed Random Sub-sampler helps to break the noise correlation of real-world and serves as an effective data augmentation for enhancing the zero-shot image denoising performance. We adopt minimum mean squared error (MMSE) estimation to mitigate the uncertainty caused by random sub-sampling and further enhance the denoising performance. With the distinctive integration of SR training strategy, Random sub-sampling and MMSE estimation, Noise2SR exhibits significant performance for single image denoising, particularly in scenarios with extremely low signal-to-noise ratios (SNR) such as HREM.
We conduct a series of experiments on both simulated and real HREM images to demonstrate the effectiveness and superiority of the proposed Noise2SR framework for HREM image denoising. The main contributions of this work can be summarized as follows:
@@ II-A Image Denoising Without Clean Signal Prior
Noise2Noise (N2N) [25] is the first work that proposes training a deep denoiser using paired noisy images of the same scene and demonstrates that it is statistically equivalent to supervised learning. Subsequently, self-supervised denoising methods have been proposed, enabling the denoiser to be trained from individual noisy images without paired noisy images. Noise2Void (N2V) [12] and Noise2Self (N2S) [11] design the masked-based blind-spot network (BSN) for self-supervised learning. Specifically, the masked-based BSN replaces certain pixels of input noisy images and predicts their value based on neighboring pixels. However, masked-based BSN suffers from the limited pixels of supervision and imperfect replacement strategies, leading to inefficient training and artifacts in the denoised results [13, 16]. To address these issues, the following works [13, 20, 14, 17, 26] have proposed to design novel BSN architectures by incorporating centrally masked convolution and dilated convolutions. However, these methods are limited by the large amount of calculation and the inflexible network structure. Different from BSN methods, another category of self-supervised learning approaches proposed generating approximate paired noisy images from individual noisy images. For example, Noisier2Noise [18] generates paired noisy images by adding additional noise to noisy images, and similar ideas are explored in [27, 23]. However, these methods require a known noise model of original noisy images to guarantee denoising performance. Additionally, Neighbor2Neighbor (NB2NB) [19] proposed to sub-sample the neighbor pixels of the noisy image to generate paired noisy images. Since a gap exists between the underlying ground truth of paired neighbor sub-sampled images, NB2NB has limited performance in low-SNR scenarios.
II-B Advances in Unsupervised and Zero-shot Image Restoration
Deep image prior (DIP) [28] first demonstrated that convolutional neural networks (CNNs) inherently learn image priors with their inductive bias. DIP proposed using early-stopping strategies for recovering corrupted images with CNNs. Subsequently, several works leverage various approaches such as denoising [29], GANs [30] or diffusion models [31] to learn the prior image distribution for zero-shot image restoration. However, these approaches require a substantial amount of data for pre-training to learn the data distribution, which is impractical in the context of HREM. Therefore, we focus on zero-shot learning restoration methods, training with only a single corrupted image without additional external data. Compared with dataset-based training, zero-shot learning is significantly more challenging to avoid overfitting. In image dehazing, several works [32, 33] have achieved remarkable performance by training on a single hazy image. For zero-shot image denoising, Self2Self [15] incorporates a dropout regularization with a blind-spot networks framework to mitigate overfitting. FBI-Denoiser [17] showed superior performance by carefully designing the BSN. ZS-Noise2Noise [22] introduced a lightweight network to achieve comparable performance in zero-shot denoising. Chen et al. [34] proposed a zero-shot method for medical image artifact reduction.
@@ II-C Electron Microscopy (EM) Denoising
Conventional spatial filter based methods have been applied to EM, such as Bilateral filter, Non-local Means, BM3D, etc. Recent, deep learning based methods have been applied in EM imaging[35, 36]. Refs. [37, 38] proposed using Cycle-GAN for STEM images denoising without paired training images. Chong et al. [39] utilized paired noisy images achieving real-world optical and electron microscopy data denoising. Refs. [40] [41] proposed the simulation-based denoising (SBD) framework creating large simulated datasets to train CNNs for denoising in a supervised learning manner. Mohan et al. [41] introduced simulated paired clean/noisy HREM images of the same substance for various imaging parameters. However, creating large simulated datasets can be a time-consuming process, and its performance will degrade sharply due to the domain gap between simulated and real noisy images. Thus, it is imperative to propose a zero-shot image denoising method that is robust to ultra-low SNR to enhance HREM data quality effectively.
@@ III Methodology
Figure 2: The figure illustrates the random sub-sampling operations in the Random Sub-sampler with sampling stride $s$. The input image is pixel first unshuffled with a stride of $s$. At each location in the pixel unshuffled image, the Random Sub-sampler randomly selects $1$ element along the channel dimension to compose the sub-sampled image $\mathbf{y}_{J^{\complement}}$. Meanwhile, the sampler sets the unsampled pixel mask $\mathbf{m}_{J}$ to 0 at the corresponding location if the element has been selected. Otherwise, it assigns a value of 1 (where black represents 0, and white represents 1). The specific example in the figure demonstrates the sub-sampling process of the Random Sub-sampler with sampling stride $(s=2)$ applied to an input image of size $4\times 4$.
@@ III-C Generating Sub-sampled Image Randomly
The stride $s$ determines the relatively sub-sampling interval and sampling ratio of noisy pixels, which have an impact on the training of Noise2SR. In Section IV-F, we conduct a comprehensive study to evaluate the influence of the sampling stride $s$ of the Random Sub-sampler. We also compare the performance between fix-location and random-location sub-sampling strategies.
@@ III-D Optimizing Network
Based on the proof provided in Section III-B, we optimize the network by minimizing the loss function that compares the unsampled pixels of the original noisy image denotes as $\mathbf{y}_{J}$, with corresponding pixels in the network output, denotes as $f_{\theta}(\mathbf{y}_{J^{\complement}})_{J}$. To facilitate the selection of these pixels, we can utilize a binary mask $\mathbf{m}_{J}$. Specifically, the unsampled pixels of the original image can be obtained by element-wise multiplication with the mask, i.e., $\mathbf{y}_{J}=\mathbf{y}\odot\mathbf{m}_{J}$. Similarly, the corresponding pixels in the network output can be obtained as $f_{\theta}(\mathbf{y}_{J^{\complement}})_{J}=f_{\theta}(\mathbf{y}_{J^{\complement}})\odot\mathbf{m}_{J}$. Here, the symbol $\odot$ represents the Hadamard product. Thus, the Noise2SR network can learn a denoising and SR function by minimizing the self-supervised loss
@@
where $\mathcal{J}$ represents the set of partitions used during the training stage, and $N$ denotes the number of images in the dataset.
@@ III-E Clean Image Restoration with MMSE Estimation
where $J_{m}^{c}$ denotes the sub-sampling partition at $m$ time.
@@ III-F1 U-Net based Encoder
We adopt the same U-Net architecture as encoder [25] with modifying the last convolution block to generate a feature map with 128 channels for Super-resolved Decoder.
@@ III-F2 Super-resolved Decoder
The Super-resolved Decoder takes the feature map from the encoder as input and generates the denoised image prediction. It is comprised of an Upsampling layer followed by three $1\times 1$ convolution layers. In this work, we employ pixel-shuffling [43] as the super-resolution strategy for the Upsampling layer. In Section IV-F, we conduct an ablation study to evaluate the impact of different super-resolution strategies on denoising performance.
@@ IV Experiments and Results
In this section, we aim to demonstrate the feasibility and effectiveness of the proposed Noise2SR framework. First, we conduct an experiment to evaluate the influence of sub-sampling on clean/noisy natural and HREM images. Then, We conduct two noise removal experiments comparing Noise2SR with eight other denoising methods. 1) Synthetic Poisson-Gaussian noise removal on simulated TEM datasets, and 2) Real noise removal on Scanning Transmission Electron Microscopy (STEM) imaging data. Moreover, we perform a comprehensive ablation study to analyze the impact of various factors, such as sub-sampling patterns, the number of samples used for MMSE estimation of the clean signal and the up-sampling strategy adopted in Super-resolved decoder.
@@ IV-A2 Implementation Details
For the training of the Noise2SR model, each iteration begins by randomly cropping a patch image of size $256\times 256$ from the original image. Subsequently, we utilize the Random Sub-sampler with sampling stride $(s=2)$ to generate a sub-sampled noisy image, which serves as the input for the network. To optimize the model, we employ the Adam optimizer [46] with the following hyper-parameters: $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. The initial learning rate is set to $10^{-4}$. We set the batch size of 12 and the training process consists of 1500 epochs. During the inference stage, we sub-sampled 50 times to generate an approximate MMSE denoised result.
@@ IV-B2 Noise simulation
$(\mathbf{x}+\mathbf{n}_{p}(\mathbf{x}))\sim{a}\mathcal{P}(\frac{1}{a}\mathbf{x}),$
@@
Figure 7: Comparison of other image denoising methods with Noise2SR on real STEM data. The top two rows display the denoising results of ZSM-5 zeolite, while the last two rows showcase the denoising results of MOR zeolite. The red and yellow arrows point to Silicon and Oxygen atoms, respectively. * indicates that the dataset-based self-supervised denoising method was performed in a zero-shot learning manner.
IV-C Comparisons on Simulated TEM Datasets

### v2

@@ I Introduction
High-resolution electron microscopy (HREM) imaging [1, 2, 3] is an indispensable tool in the fields of materials science and nanotechnology. HREM enables direct visualization of structures at the atomic level through interactions between the sample and high-energy electrons. However, HREM is inevitably susceptible to noise due to inherent properties of electron beams and detection process etc. For example, the low-dose conditions are often applied to imaging electron-beam sensitive materials to minimize the damage from the electrons [4, 5]. Capuring the dynamic events at kilohertz frame rates with direct electronic detection systems [6, 7], the image is affected by severe shot noise due to the shortened exposure time. Improving signal-to-noise (SNR) is critical for HREM to enhance image quality and facilitate accurate information extraction.
@@
Recently, data-driven methods based on deep learning (DL) [8, 9, 10] have obtained favorable performance compared to conventional methods in image denoising. However, applying supervised DL denoising methods to electron microscopy (EM) images is challenging due to the lack of paired noisy-clean image datasets. In the absence of ground-truth images, several self-supervised image denoising methods[11, 12, 13, 14, 15, 16, 17, 18, 19, 20] have been proposed. Some works [12, 11, 13, 17] utilize blind-spot networks (BSNs) to prevent identical mappings in self-supervised learning. BSNs predict the clean value of a noisy pixel without using the information from that pixel itself. To build BSN, Noise2Void [12] and Noise2Self [11] initially employ a masked-based strategy, while subsequent works [14, 13, 21, 17] propose tailored network designs. Huang et al. proposes Neighbor2Neighbor (NB2NB) [19] to generate paired noisy images from subsampling paired neighbor pixels for self-supervised training. While NB2NB and its variants [22, 23] fail to address the intensity gap issue between neighbor pixels, resulting in relatively poor performance in low-SNR scenarios. Additionally, some works [18, 24] introduce explicit noise modeling to generate training paired noisy images by adding synthetic noise. However, in HREM imaging, the noise distribution of real HREM is unknown and hard to estimate. Furthermore, the scarcity of HREM data poses a significant challenge for these methods since their performance tends to degrade with insufficient data training [15, 23].
@@
In this paper, we propose an efficient zero-shot self-supervised denoising framework for HREM images named Noise2SR. We propose to train a denoising network with paired noisy images with different resolutions, which is conducted via super-resolution (SR) strategy. Inspired by NB2NB [19], we introduce the Random Sub-sampler module to generate sub-sampled noisy images that form a noisy pair with the original noisy image. Unlike NB2NB, we utilize paired noisy images with different resolutions for training. We indicated that the sub-sampled noisy image and the original noisy image have a consistent underlying clean image. Meanwhile, we provide theoretical proof that the Noise2SR training scheme is statically equivalent to using a clean image for supervision. Combined with the SR-based training strategy, we address the coordinate mismatch and intensity gap issues present in paired neighbor pixels in NB2NB. The proposed Random Sub-sampler helps to break the noise correlation of real-world and serves as an effective data augmentation for enhancing the zero-shot image denoising performance. We adopt minimum mean squared error (MMSE) estimation to mitigate the uncertainty caused by random sub-sampling and further enhance the denoising performance. With the distinctive integration of SR training strategy, Random sub-sampling and MMSE estimation, Noise2SR exhibits significant performance for single image denoising, particularly in scenarios with extremely low signal-to-noise ratios (SNR) such as HREM. We conduct a series of experiments on both simulated and real HREM images to demonstrate the effectiveness and superiority of the proposed Noise2SR framework for HREM image denoising. The main contributions of this work can be summarized as follows:
@@ II-A Image Denoising Without Clean Signal Prior
Noise2Noise (N2N) [26] is the first work that proposes training a deep denoiser using paired noisy images of the same scene and demonstrates that it is statistically equivalent to supervised learning. Subsequently, self-supervised denoising methods have been proposed, enabling the denoiser to be trained from individual noisy images without paired noisy images. Noise2Void (N2V) [12] and Noise2Self (N2S) [11] design the masked-based blind-spot network (BSN) for self-supervised learning. Specifically, the masked-based BSN replaces certain pixels of input noisy images and predicts their value based on neighboring pixels. However, the replacement strategy cannot entirely prevent the network from learning identical mappings in self-supervised denoising, as pointed out by [16]. Additionally, the masked-based BSNs suffer from limited pixels for supervision, resulting in inefficient training [13] and artifacts in the denoised results. To address these issues, the following works [13, 21, 14, 17, 27] have proposed to design novel BSN architectures by incorporating centrally masked convolution and dilated convolutions. However, these methods are limited by the large amount of calculation and the inflexible network structure. Different from BSN methods, another category of self-supervised learning approaches proposed generating approximate paired noisy images from individual noisy images. For example, Noisier2Noise [18] generates paired noisy images by adding additional noise to noisy images, and similar ideas are explored in [28, 24]. However, these methods require a known noise model of original noisy images to guarantee denoising performance. Additionally, Neighbor2Neighbor (NB2NB) [19] proposed to sub-sample the neighbor pixels of the noisy image to generate paired noisy images. Since a gap exists between the underlying ground truth of paired neighbor sub-sampled images, NB2NB has limited performance in low-SNR scenarios.
II-B Advances in Zero-shot Image Denoising
Zero-shot learning in image denoising refers to training a denoiser using only a single noisy image without any external data. This approach is particularly valuable in scenarios where training data is scarce, such as in HREM. The main challenge is avoiding overfitting, which can significantly reduce the effectiveness of the aforementioned self-supervised denoising methods as training data decreases. Deep image prior (DIP) pioneered zero-shot image denoising by utilizing the early-stopping strategies in convolutional neural networks (CNNs) training. With the prosperity of self-supervised denoising, Self2Self (S2S) trained a zero-shot denoiser that incorporates a dropout regularization with a blind-spot networks framework to mitigate overfitting. FBI-Denoiser [17] showed superior performance of zero-shot denoising in Poission-Gaussian noise denoising through careful BSN design. Recent studies [22, 23] explored fast and efficient zero-shot denoisers by introducing a novel sub-sampling strategy or a lightweight network.
@@ II-C Electron Microscopy (EM) Denoising
Conventional spatial filter based methods have been applied to EM, such as Bilateral filter [29, 30], Non-local Means [31, 32], BM3D[33], etc. Recent, deep learning based methods have been applied in EM imaging [34, 35]. Refs. [36, 37] proposed using Cycle-GAN for STEM images denoising without paired training images. Chong et al. [38] utilized paired noisy images achieving real-world optical and electron microscopy data denoising. Refs. [39] [40] proposed the simulation-based denoising (SBD) framework creating large simulated datasets to train CNNs for denoising in a supervised learning manner. Mohan et al. [40] introduced simulated paired clean/noisy HREM images of the same substance for various imaging parameters. However, creating large simulated datasets can be a time-consuming process, and its performance will degrade sharply due to the domain gap between simulated and real noisy images. Thus, it is imperative to propose a zero-shot image denoising method that is robust to ultra-low SNR to enhance HREM data quality effectively.
@@ III Methodology
Figure 2: Illustration of the operations in the Random Sub-sampler with sampling stride $s$ to generate sub-sampled image $\mathbf{y}_{J^{\complement}}$ and unsampled mask $\mathbf{m}_{J}$. First, the input image is applied pixel unshuffling with a stride of $s$. At each location in the pixel unshuffled image, the Random Sub-sampler randomly selects $1$ element along the channel dimension to compose the sub-sampled image $\mathbf{y}_{J^{\complement}}$. Meanwhile, the sampler sets the unsampled pixel mask $\mathbf{m}_{J}$ to 0 at the corresponding location if the element has been selected. Otherwise, it assigns a value of 1 (where black represents 0, and white represents 1). The specific example in the figure demonstrates the sub-sampling process of the Random Sub-sampler with sampling stride $(s=2)$ applied to an input image of size $4\times 4$.
@@ III-C Generating Sub-sampled Image Randomly
In section IV-F, we evaluate the effectiveness of randomness in sub-sampler on the performance. The stride $s$ determines the sub-sampling interval, which helps to break the spatial correlation of noise. It also controls the sampling ratio of noisy pixels and affects the relative receptive field of the network, both of which influence the training process of Noise2SR. In Section IV-G, we conduct a comprehensive study to evaluate the influence of the sampling stride $s$ of the Random Sub-sampler. Recent studies [19, 20] adopt similar sub-sampling strategies. However, our Random Sub-sampler generates paired noisy images at different resolutions while their sub-sampled images are at the same resolution.
@@ III-D Optimizing Network
Based on the proof provided in Section III-B, we optimize the network by minimizing the loss function that compares the unsampled pixels of the original noisy image denotes as $\mathbf{y}_{J}$, with corresponding pixels in the network output, denotes as $f_{\theta}(\mathbf{y}_{J^{\complement}})_{J}$. To facilitate the selection of these pixels, we can utilize a binary mask $\mathbf{m}_{J}$. Specifically, the unsampled pixels of the original image can be obtained by element-wise multiplication with the mask, i.e., $\mathbf{y}_{J}=\mathbf{y}\odot\mathbf{m}_{J}$. Similarly, the corresponding pixels in the network output can be obtained as $f_{\theta}(\mathbf{y}_{J^{\complement}})_{J}=f_{\theta}(\mathbf{y}_{J^{\complement}})\odot\mathbf{m}_{J}$. Here, the symbol $\odot$ represents the Hadamard product. Thus, the Noise2SR network can learn a denoising and SR function by minimizing the self-supervised loss:
@@
where $\mathcal{J}$ represents the set of partitions used during the training stage.
@@ III-E Clean Image Restoration with MMSE Estimation
where $J_{m}^{\complement}$ denotes the sub-sampling partition at $m$ time.
@@ III-F1 U-Net based Encoder
We adopt the same U-Net architecture as encoder [26] while modifying the last convolution block to generate a feature vector $\mathbf{v}$ with 128 channels for the Super-resolved decoder.
@@ III-F2 Super-resolved Decoder
The decoder takes the feature vector as input to generate the denoised image $\hat{\mathbf{x}}$. It consists of an upsampling layer followed by three convolutional layers, each with a 1$\times$1 kernel size. In this work, we employ pixel-shuffling [42] as the super-resolution strategy for the upsampling layer. In Section IV-G, we conduct an ablation study to evaluate the impact of different super-resolution strategies on denoising performance.
@@ IV Experiments and Results
In this section, we aim to demonstrate the feasibility and effectiveness of the proposed Noise2SR framework. First, we evaluate the performance of Noise2SR compared with eight other denoising methods on simulated and real noise removal in HREM data. Then, we conduct comprehensive ablation studies to analyze the impact of key components, including the Random Sub-sampler, MMSE estimation of the clean signal, and the architecture of the encoder and decoder in the Noise2SR network. Finally, we demonstrate some failure cases to demonstrate the extreme performance limits of Noise2SR and discuss the potential approaches for improvement.
@@ IV-A2 Implementation Details
For the training of the Noise2SR model, each iteration begins by randomly cropping a patch image of size 128 $\times$ 128 from the original image. Subsequently, we utilize the Random Sub-sampler with sampling stride $(s=2)$ to generate a sub-sampled noisy image, which serves as the input for the network. To optimize the model, we employ the Adam optimizer [44] with the following hyper-parameters: $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. The initial learning rate is set to $10^{-4}$. We set the batch size of 12 and the training process consists of 1500 epochs. During the inference stage, we sub-sampled 50 times to generate an approximate MMSE denoised result.
@@ IV-B2 Noise simulation
$(\mathbf{x}+\mathbf{n}_{p}(\mathbf{x}))\sim{a}\mathcal{P}\left(\frac{1}{a}\mathbf{x}\right),$
@@
IV-C Comparisons on Simulated Noisy TEM Datasets
@@ IV-D1 Low-Dose Te Crystal STEM Data
Figure 7: Comparison of other image denoising methods with Noise2SR on real STEM data. The top two rows display the denoising results of ZSM-5 zeolite, while the last two rows showcase the denoising results of MOR zeolite. The orange and blue arrows point to Silicon and Oxygen atoms, respectively. * indicates that the dataset-based self-supervised denoising method was performed in a zero-shot learning manner.
IV-D2 ZSM-5 & MOR Zeolites STEM Data
To further apply this method, two kinds of zeolites, ZSM-5 zeolite and MOR zeolite were imaged under very low-dose conditions and then denoised. It is worth mentioning that zeolites, as one kind of import nano-porous materials that were widely used in catalysis, separation, etc, are extremely sensitive to electron beams. Thus, low electron dose conditions are applied, resulting in low SNR of STEM images which hinder the visualization of atomic structure.
Figure 8: Training time of denoiser with zero-shot learning v.s the PSNR results on simulated TEM corrupted by Poisson-Gaussian noise with $(a=0.05,b=0.02)$.
IV-E Comparison of Time Required for Zero-shot Denoiser
A zero-shot denoiser is an image-specific denoiser that requires training on the target noisy image. However, the excessive training time for zero-shot denoisers can limit their practical applications. In our experiments, we found that the training time of zero-shot denoiser takes several minutes, while the inference time is seconds, which can be neglected. We compared the training time required by different denoisers to achieve the performance results shown in Table I. Additionally, we compared the performance of Noise2SR with early stopping, which takes approximately the same training time as Noise2Self. To better illustrate the balance between training time and performance, we included Fig. 8. The results show that, compared to FBI-D, Noise2SR can achieve similar performance in less time. Furthermore, compared to Noise2Self, Noise2SR delivers better performance within the same training time. By extending the training time, Noise2SR can significantly outperform other methods.